2x(2) - 3x - 5 = 0
1
Expert's answer
2013-04-10T07:39:56-0400
2x^2 - 3x - 5 = 0
We're dealing with a quadratic equation. Let's find its discriminant:
D=9+4*5*2=49>0 therefore given quadratic equation has two distinct real roots. They are
x1,2=(3+-sqrt(D))/4=(3+-7)/4
x1=10/4-5/2, x2=-1
So the answer is x=5/2; x=-1.